18.
Find the slope of the line passing through the points (15, 6) and (10, 1).



Answer by venugopalramana(3286)

Find the slope of the line passing through the points (15, 6) and (10, 1).

SLOPE =DIFFERENCE IN Y COORDINATES/DIFFERENCE IN X COORDINATES
=(6-1)/(15-10)=5/5=1